assume that x has a normal distribution, and find the indicated probablility. The mean is 60.0 and the standard deviation is 4.0. Find the probability that x is less than 53.0
Answer by stanbon(75887)

assume that x has a normal distribution, and find the indicated probablility. The mean is 60.0 and the standard deviation is 4.0. Find the probability that x is less than 53.0
--------
z(53) = (53-60)/4.0 = -7/4
-----
P(x < 53) = P(z < -7/4) = normalcdf(-100,-7/4) = 0.0401
======================
Cheers,
Stan H.